OpenStudy (anonymous):

Graph the Following System of Equations y=x+2 y=-x+2 Not quite sure how to do this

OpenStudy (australopithecus):

did you just reask this question? wow

OpenStudy (australopithecus):

since they both equal y sub one into the other x + 2 =-x + 2 solve for x

OpenStudy (australopithecus):

so 2x = 0 x = 0 so y = 0 + 2 y = 2

OpenStudy (australopithecus):

if I sub x=0 into the other equation I should get the same thing y = 2 - x y = 2 - 0 y = 2

OpenStudy (australopithecus):

thus the two equations intersect at (0,2)

OpenStudy (australopithecus):

if you dont understand this ask questions dont just keep asking the same questions over and over

OpenStudy (anonymous):

I diont remember posting this twice, if i did it was an accident but thankyou very much for your help! @Australopithecus